OpenStudy (anonymous):

standard form of the line which passes through (6, 3) and has a slope of two thirds?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(6\quad ,&3) \end{array} \\\quad \\ slope = m= \cfrac{2}{3} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

y-y1 = m(x-x1) m is the slope plug in the points then turn it into a standard form of ax+by = c

OpenStudy (anonymous):

Alright thank you! I was just double checking my answer! :)

OpenStudy (anonymous):

Does anyone know how to find this: Given f(x) = the quantity of 5x plus 1, divided by 2 solve for f-1(8). I'm not very sure of my process. Thanks!

OpenStudy (jdoe0001):

\(\bf f(x)=y=\cfrac{5x+1}{2}\qquad inverse\implies x=\cfrac{5y+1}{2}\Leftarrow f^{-1}\) notice, all we did was swap the 2 variables about then solve for "y"

OpenStudy (jdoe0001):

once you've solve for "y", that will be \(\bf f^{-1}\) and then just pass the "8" to it like \(\bf f^{-1}(8)\)

OpenStudy (anonymous):

Yes, I've got that, I'm just not sure what to do with the 8

OpenStudy (jdoe0001):

\(\bf f(x)=y=\cfrac{5x+1}{2}\qquad inverse\implies x=\cfrac{5y+1}{2}\\ \quad \\ f^{-1}=y=\cfrac{2x-1}{5}\qquad f^{-1}(\color{red}{8})=\cfrac{2(\color{red}{8})-1}{5}\)

OpenStudy (anonymous):

Ohhh, i see I see! Thanks! :)

OpenStudy (jdoe0001):

yw